A computer can perform 30 identical tasks in 6 hours. At that rate, What is the minimum number of computers that should be assigned to complete 80 tasks within 3 hours.
Thanks.

Answer by Schaman_Dempster(26)

A computer can perform 30 identical tasks in 6 hours.
Rate at which a computer works = number of tasks/ number of hours = 30/6 = 5 tasks per hour
We can say that a computer performs 5 tasks in an hour.
So, a computer performs 3*5 = 15 tasks in 3 hours.
Now, if x number of computers are operated for 3 hours. Then they can do number of tasks = number of tasks done by a computer in 3 hours times the number of computers
Therefore, Number of tasks done by x computers in 3 hours = 15*x = 15x
Now, this number of tasks should equal 80 as mentioned in the problem.
Therefore, 15x = 80
x = 80/15
x = 16/3
x = 5 and 1/3
Oh we can't have fraction number of computers.
Hence, we can say to complete 80 tasks, there should be 6 computers operating simultaneously.